HelpMePlz:

Given the following triangle, solve for the missing side: 9.32 8.67 7.34 6.4

HelpMePlz:

http://prnt.sc/d5xsoe

Champion:

Apply Pythagoras Theorem, Hypotenuse ^2 = c^2 = a^2+b^2 Here the two known sides are a=4, b=5. So calculate c (hypotenuse) using the above formula. Use your calculator, and don't forget to calculate the square-root of c^2.